I have managed to work the rest of this question but the last bits are hell.Please help.
I am trying to find the GRADIENT of the line PERPENDICULAR to AB
(i)
A(3,2)................B(4,-1)
(ii)
A(-6,3)...............B(6,3)

Answer by venugopalramana(3286)

gradient of line joining (x1,x2),(y1,y2)is given by (y2-y1)/(x2-x1)
product gradients of 2 perpendicular lines =g1*g2 = -1
so we have gradient of ab =(-1-2)/(4-3)=-3/1=-3=g1
g2=-1/g1=-1/-3=1/3.............the other example also is similar
A(-6,3)...............B(6,3)
g1=(3-3)/(6+6)=0=g1
g2=-1/0= -infinity